Grade 12th passAlgebra

The rational numbers that are equal to their reciprocals.

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

The rational numbers that are equal to their reciprocals are 1 and -1.
Reason:
1=1/1
Reciprocal of 1 = 1/1 = 1
Similarly, Reciprocal of -1 = - 1

Thanks